OpenStudy (anonymous):

Zartan's vine, initially 6m long has a Young's modulus of .1 GN/m^2. If zartans weight causes a stress of 2.5*10^5 N/m^2, then A)Determine the vnes change in length B)the new length of the rope

OpenStudy (anonymous):

\[l_{total} =l_o+\Delta l\]

OpenStudy (anonymous):

\[Y=\frac{a}{\epsilon}=\frac{stress}{strain}=\frac{\frac{F}{A}}{\frac{\Delta l}{l_0}}\] \[=\frac{Fl_0}{\Delta l A}\]

OpenStudy (anonymous):

so now we have \[Stress*\frac{\Delta L}{l_0}=Y\] \[\frac{Y}{stress}=\frac{\Delta l}{l_0}\] \[\frac{Yl_0}{stress}=\Delta l\]

OpenStudy (anonymous):

\[Stress*\frac{l_0}{\Delta l}=Y\] \[\frac{Stress*l_0}{Y}=\Delta l\] \[\frac{2500000*6}{100000000}=.15m=15cm\] \[l_{total}=l_0+\Delta l=6+.15=6.15m\]